# Reasoning Language Models: A Blueprint

### v1

@@ 7.5.1 Example Results
We show a visualization of the token probabilities in Figure 8 for the following model output shown in Figure 7.
Figure 7: Example model output with highlighted tokens. The output has been colored in the following way: Purple if the highest probability is below 0.8 (not so certain but no contention), blue if the second highest probability is above 0.1 (very certain, but maybe another one) and red if both are true (uncertain).
@@
Figure 8: Token probabilities of example model outputs We show the two highest probabilities as well as the sum of the other probabilities.
By examining the highest probabilities, the second-highest probabilities, and the sum of the remaining ones, one can get a first impression of the underlying distribution, that can later be quantified through metrics. There are regions, where the top two probabilities are very close and all other values are significantly smaller. This scenario likely represents a situation where forking in the sampling process could disproportionately benefit future outcomes. Other potential key regions include instances where the model is highly confident (with a high first probability, a low second, and minimal remaining probabilities) and cases where the remaining probabilities are nearly as high as the first and second, indicating a flatter distribution and high uncertainty. In such cases, rather than forking, it may be more beneficial to correct potential errors in previous generations.
@@
Figure 9: Uncertainty metrics (variance, entropy, VarEntropy, and the Gini coefficient) plotted against the output token sequence.
We now further investigate these results in Figure 9 with the help of the uncertainity metrics introduced earlier. It is evident that the zero-shot prompt seems less uncertain and therefore probably provides better output. However the use of the value 472 as a token could mean that the question has been incorporated at least into the tokenizer and may also have been leaked into the training in some form. Another immediate standout is the high uncertainty associated with all scores around the <thought> tokens, which seems challenging to predict. Additionally, the token granularity results in most words being represented by just a single token, which implies that, statistically, the next token (given a random one) is likely to be roughly evenly distributed across the English language. A low uncertainty indicates the model has high confidence in predicting the next token based on the given context. Consequently, these uncertainty values reflect the model’s learned statistical patterns rather than the actual statistical structure of the language used.
Finally we illustrate ways to harness these metrics, specifically a combination of entropy and VarEntropy, in the design of a reasoning strategy. Cases of a high VarEntropy and a lower Entropy could benefit from a tree search, since some of the outcomes have a equally highly probability, whereas the other outcomes have a low probability. In contrast, a low VarEntropy and a high Entropy indicates situations that could benefit from a clarification as all outcomes are equally improbable. We observe that these metrics highlight tokens a lot of times which are not essential for the overall task but for design and formatting decisions. The Gini coefficient shows more promising patterns in vital areas.
@@ 8 Example Insights for Effective RLMs
9 Benchmarking
Sun et al. [85] provide a clear distinction between various types of reasoning including mathematical, logical, casual, and commonsense. Below, we highlight a selection of benchmarks for each category.
@@ 9.4 Commonsense Reasoning
9.5 Benchmarking RLMs
Including multiple models in a reasoning scheme tends to increase its output variance. Therefore, we emphasize the importance of benchmarking RLMs on sufficiently large sample sizes to ensure fair comparisons across different approaches. Benchmarks such as AIME or AMC, which only provide a limited number of samples in the two-digit range, pose a risk of selective reporting, where researchers might highlight subsets of results where their models performed well.
Experimental results (see Figure 10) testing the variability of LLM generation indicate that achieving low error variability in the single-digit percentage range requires evaluating on over 500 samples. As RLMs can have much larger variability due to the increased complexity and based on these findings, we recommend that individual benchmarks contain at least 200 samples each and that a minimum of 500 samples be evaluated per category to ensure robust performance comparisons.
@@ 11 Conclusion
This work introduces a comprehensive blueprint for reasoning language models (RLMs), providing a flexible and modular toolbox that clarifies the intricate design and operation of these advanced systems. By encompassing diverse reasoning structures, operations, and training schemes, the blueprint serves as a robust foundation for constructing, analyzing, and extending RLMs tailored to various applications. The x1 implementation further enhances this contribution, offering a modular, minimalist, and user-friendly platform for experimentation and the rapid prototyping of novel RLM architectures.
Our blueprint and x1 pave the way for future advancements in reasoning AI. The blueprint can inspire the development of new RLM architectures, such as nested designs where reasoning structures like trees and graphs are embedded hierarchically for solving multi-layered tasks. x1 serves as a playground for exploring these novel models, enabling researchers to experiment with enhancements such as refined training paradigms, innovative reward structures, and adaptive search strategies.
@@ Acknowledgements
We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault, Piz Daint and Alps machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.
@@ Appendix A Mathematical Foundation of Markov Decision Processes for Reasoning Tasks
TABLE II: Notation Overview
@@
$r(s,a,s^{\prime})$ Reward received after taking action $a$ in state $s$ resulting in state s’.
@@
In this section, we provide a rigorous mathematical framework for RLMs by integrating Markov Decision Processes (MDPs) with Monte Carlo Tree Search (MCTS) algorithm. The MDP serves as a foundational formulation for modeling reasoning chains, which constitute the reasoning structure of the RLMs. Simultaneously, MCTS serves as an efficient search algorithm for exploring and navigating the expansive space of possible reasoning chains. The resulting state space is then used as a basis for modeling the RLM. An overview of the notation used in this section is provided in Table II.
@@ A.1 Markov Decision Process
Solving an MDP
A trajectory $\tau_{\pi}=(s_{0},a_{0},\ldots,s_{T},a_{T},s_{T+1})$ is a sequence of interleaved states and actions, selected according to the policy $\pi$. Each trajectory starts at an initial state $s_{0}\in\mathcal{S}$ and ends with $s_{T+1}\in\mathcal{S}$ which represents the terminal state where no further actions can be taken.
A policy $\pi(a\mid s)$ is a function assigning a probability distribution over the action space to a given state $s$; $\pi:\mathcal{S}\rightarrow\Delta(A)$. The expression $\pi(a\mid s)$ denotes the probability of selecting the action $a$ in the state $s$ according to the policy $\pi$.
State value function $V(s_{t})$ represents the expected cumulative future reward for a given state $s_{t}$ under policy $\pi$:
@@
The State-Action value function $Q(s_{t},a_{t})$ The state-action value function $Q(s_{t},a_{t})$ extends the state value function to evaluate a specific action $a_{t}$:
@@
$\stackrel{{\scriptstyle}}{{=}}r(s_{t},a_{t})+\gamma\mathbb{E}_{s_{t+1}}\left[V_{\pi}(s_{t+1})\mid s_{t},a_{t}\right]$
, where Bellman’s equation is used in the second equality.
A.1.1 MDPs in the RLM Setting
In the context of RLMs a state $s\in\mathcal{S}$ is typically defined as a sequence of reasoning steps $s=(z_{0}\dots z_{n})$, where each reasoning step $z_{i}$ is a sequence of $M_{i}$ tokens $z_{i}=(t_{i}^{0},\dots,t_{i}^{M_{i}})$. Each $t_{i}^{j}$ is a token from the RLM’s vocabulary, and the total number of tokens per reasoning step $M_{i}$ can vary. One can use a special token $t_{M_{i}}=t_{end}$ to indicate the end of the reasoning step. Typically, the initial query $q$ is used as the first reasoning step $z_{0}=q$. In the study of RLMs, an action $a\in\mathcal{A}_{s}$ usually represents appending a new reasoning step $z^{(a)}$ to the current state $s=(z_{0},...,z_{n})$ resulting in a new state $s^{\prime}=(z_{0},...,z_{n},z^{(a)})$. Since every action is uniquely associated with exactly one reasoning step, the transitions between states are deterministic i.e for $s=(z_{0},...,z_{n})$ and $s=(z_{0},...,z_{n},z_{n+1})$.
$p(s,a,s^{\prime})=\begin{rcases}\begin{dcases}1&\text{ if }z_{n+1}=z^{(a)}\\ 0&\text{ if }z_{n+1}\neq z^{(a)}\\ \end{dcases}\end{rcases}$
The definition of the reward function depends on the specific task. A reward commonly seen in reasoning tasks assigns non-zero reward only in the terminal states and hence only at the final reasoning step. This approach reflects the fact that for most tasks, the only final answer can be evaluated against the ground-truth solution to the original query. We call such reward functions sparse to cleary distinguish it from other setting in which an intermediate rewards can be observed by the algorithm in the non-terminal states. The discount factor $\gamma$ determines how future rewards influence the current decision-making process. A higher discount factor ($\gamma\rightarrow 1$) places greater emphasis on long-term reasoning success, allowing the model to generate long reasoning sequences, while a lower discount factor prioritizes immediate rewards, incentivizing faster progress and shorter reasoning sequences.

### v2

@@ 7.6.3 Example Results
Figure 7 and 8 illustrate example model outputs and their respective token probability distributions. By analyzing the highest probabilities, the second-highest probabilities, and the sum of the remaining probabilities, we gain valuable insights into the underlying token distribution, which can subsequently be quantified through the uncertainty metrics discussed earlier.
[…]
(b) <thought>First, we need to find the square root of 20, so let’s use a calculator to get an approximate value of $\sqrt{20}\approx 4.472$.<thought><thought>Since we’re dealing with the ceiling function, we should consider the smallest integer greater than or equal to the square root of 20, which is the ceiling of $\sqrt{20}$. </thought><thought>We can also consider the perfect squares that are closest to 20, such as 16 and 25, to estimate the value of $\sqrt{20}$ and then apply the ceiling function.</thought>
(c) <thought>First, we need to find the square root of 20, which is approximately 4.472, so we can round up to the nearest whole number using the ceiling function, which gives us 5.</thought><thought>We can evaluate the expression $\lceil{\sqrt{20}}\rceil^{2}$ by first finding the square root of 20, then rounding up to the nearest whole number, and finally squaring the result.</thought><thought>Since $\lceil{x}\rceil$ is the ceiling function, we can rewrite the expression as $\lceil{\sqrt{20}}\rceil^{2}=\lceil{4.472}\rceil^{2}=5^{2}=25$.</thought>
(d) <thought>First, we need to calculate the square root of 20.</thought><thought>We can start by finding the square root of 20, which is approximately 4.472, and then round up to the nearest whole number using the ceiling function.</thought><thought>Another approach is to recognize that 20 is between the perfect squares 16 and 25, so we can use this information to estimate the ceiling of the square root of 20.</thought>
@@
Figure 8: Probabilities of the first 64 tokens of example model outputs. We show the two highest probabilities as well as the sum of the other probabilities.
In Figures 8(a) and 8(d), specific regions emerge where the top two probabilities are very close, while the remaining probabilities are significantly smaller. Such regions likely indicate scenarios where forking the reasoning process (e.g., exploring multiple paths) could disproportionately benefit future outcomes, as the competing high-probability tokens suggest alternative plausible continuations. Conversely, in instances where the first probability is notably high, with much lower second and remaining probabilities, the model exhibits strong confidence in a single continuation. These cases are conducive to more deterministic reasoning, as forking may be unnecessary.
Additionally, regions with a relatively high sum of the remaining probabilities (close to the top two) highlight flatter distributions with high uncertainty. These scenarios signal a need for cautious reasoning, where clarification or additional contextual refinement may help reduce ambiguity. For instance, such uncertainty may suggest that the model has not yet committed to a specific path and could benefit from revisiting earlier reasoning steps to address potential errors or misalignments.
[…]
(b) <thought>First, we need to find the square root of 20, so let’s use a calculator to get an approximate value of $\sqrt{20}\approx 4.472$.<thought><thought>Since we’re dealing with the ceiling function, we should consider the smallest integer greater than or equal to the square root of 20, which is the ceiling of $\sqrt{20}$. </thought><thought>We can also consider the perfect squares that are closest to 20, such as 16 and 25, to estimate the value of $\sqrt{20}$ and then apply the ceiling function.</thought>
(c) <thought>First, we need to find the square root of 20, which is approximately 4.472, so we can round up to the nearest whole number using the ceiling function, which gives us 5.</thought><thought>We can evaluate the expression $\lceil{\sqrt{20}}\rceil^{2}$ by first finding the square root of 20, then rounding up to the nearest whole number, and finally squaring the result.</thought><thought>Since $\lceil{x}\rceil$ is the ceiling function, we can rewrite the expression as $\lceil{\sqrt{20}}\rceil^{2}=\lceil{4.472}\rceil^{2}=5^{2}=25$.</thought>
(d) <thought>First, we need to calculate the square root of 20.</thought><thought>We can start by finding the square root of 20, which is approximately 4.472, and then round up to the nearest whole number using the ceiling function.</thought><thought>Another approach is to recognize that 20 is between the perfect squares 16 and 25, so we can use this information to estimate the ceiling of the square root of 20.</thought>
@@
Figure 9: Uncertainty metrics (variance, entropy, VarEntropy, and the Gini coefficient) plotted against the first 64 tokens of the output token sequence.
Figure 9 further analyzes these results using metrics such as variance, entropy, VarEntropy, and the Gini coefficient. In Figure 9(a), a zero-shot prompt demonstrates lower uncertainty overall, suggesting that it yields more confident predictions and potentially higher-quality outputs. However, the presence of specific high-probability tokens (e.g., “472”) raises concerns about potential data leakage into the training set or the tokenizer, which could bias the results. Another notable observation is the high uncertainty associated with <thought>tokens, which appear challenging for the model to predict accurately. This highlights the complexity introduced by token granularity, where most words correspond to single tokens, resulting in a roughly even distribution for the next token across the vocabulary in some contexts.
The uncertainty metrics provide actionable insights for reasoning strategy design. For example, cases with high VarEntropy and low entropy indicate a distribution where a few outcomes dominate, making tree-based search strategies effective. These strategies prioritize exploring high-probability outcomes while avoiding unnecessary evaluations of less probable branches. In contrast, low VarEntropy and high entropy reflect a flat distribution where no clear outcome dominates. Such cases could benefit from clarification mechanisms or intermediate step refinements to reduce ambiguity before proceeding further.
Interestingly, the Gini coefficient often highlights critical regions more effectively than other metrics. In vital reasoning areas, it captures the inequality in token probabilities, helping to identify tokens that significantly influence the reasoning process. This contrasts with metrics like entropy and VarEntropy, which may also flag tokens related to formatting or stylistic choices, providing less task-specific utility.
Overall, these visualizations and metrics emphasize the importance of analyzing token probability distributions to design effective reasoning strategies. By leveraging the nuanced patterns revealed by these metrics, models can better adapt to uncertainty, balance exploration and exploitation, and optimize decision-making during the reasoning process.
7.7 Benchmarking RLMs
Our experience with benchmarking RLMs highlights critical considerations for ensuring fair and reliable performance comparisons. Incorporating multiple models within a reasoning scheme often increases output variance, emphasizing the need for benchmarking on sufficiently large sample sizes. Benchmarks with limited sample sizes, such as AIME or AMC, which often provide only a two-digit range of samples, risk selective reporting. This occurs when researchers focus on subsets of results where their models perform well, rather than reflecting the true variability of their systems.
Experimental findings (Figure 10) demonstrate that achieving low error variability, within a single-digit percentage range, requires evaluation across at least 500 samples. Given the inherent complexity of RLMs, which often exhibit greater variability than simpler LLM setups, these results suggest specific sample size thresholds. We recommend that individual benchmarks contain at least 200 samples per category, with a minimum of 500 samples evaluated across all categories to ensure statistically robust comparisons. Adhering to these guidelines would in many cases mitigate variability-driven biases and facilitate more transparent assessments of RLM performance across different approaches.
@@ 8 Example Insights for Effective RLMs
9 Benchmarks for RLMs
We now outline benchmarks related to RLMs. Sun et al. [135] provide a clear distinction between various types of reasoning including mathematical, logical, casual, and commonsense. Below, we highlight a selection of benchmarks for each category. We also include additional categories related to the realm of RLMs, namely, coding related benchmarks and benchmarks that involve reasoning utilities such as tools or RAG. We show the benchmarks in Figure 11.
Figure 11: Overview of benchmarks for RLMs.
@@ 9.1 Mathematical Reasoning
Other related benchmarks include MATH-401 [164], MultiArith [118], AddSub [61] CHAMP [98], MathQA [5], ARB [123], FIMO [85], Geometry3K [88], GeoQA [26], UniGeo [24], miniF2F [175], LeanDojo [159], TheoremQA-MATH [29], TRIGO [157], LISA [69], MathVista [87], ChartQA [99], TABMWP [89], MultiHiertt [173], and SCIBENCH [148].
@@ 9.2 Logical Reasoning
Other benchmarks include ProofWriter [137], FOLIO [54], WANLI [84], CLUTRR [130], Adversarial NLI [106], AbductionRules [162], and Adversarial ARCT [107].
9.3 Coding
There also exist benchmarks related to how well a given model can code. These include ODEX [151], SWE-bench [71], DS-1000 [76], APPS [57], MBPP [6], and HumanEval [27].
@@ 9.5 Commonsense Reasoning
Other related benchmarks include Social IQa [120], SWAG [165], HellaSWAG [166], CommonSenceQA [138], PIQA [19], PHYRE [7], OpenBookQA [102], CConS [74], WinoGrande [119], and FactCC [75].
9.6 Reasoning Utilities
Benchmarking capabilities of RLMs related to reasoning utilizies involve testing the capabilities of an RLM in how it acts as an agent. This includes benchmarks such as GAIA [66], WebArena [177], Mind2Web [41], WebShop [160], ALFWorld [126], AgentBench [86], AgentGym [154], and AgentBoard [21]. Another line of related benchmarks tests the RAG capabilites [25, 156, 93, 44].
@@ 11 Conclusion
This work introduces a comprehensive blueprint for reasoning language models (RLMs), providing a flexible and modular toolbox that demystifies the intricate design and operation of these advanced systems. By encompassing diverse reasoning structures, operations, and training schemes, the blueprint establishes a robust foundation for constructing, analyzing, and extending RLMs tailored to various applications. The accompanying x1 implementation enhances this contribution, offering a modular, minimalist, and user-friendly platform for experimentation and rapid prototyping of novel RLM architectures.
Our blueprint and x1 pave the way for several exciting avenues of future research and development in reasoning AI. One example is Trace-Based Supervision (TBS), which extends process-based supervision by incorporating labeled traces of traversal through reasoning structures. TBS has the potential to train more powerful implicit RLMs capable of internalizing reasoning structures and improving generalization.
The work also explores new directions in value and reward modeling, introducing a hierarchy of models and formally identifying several recent designs as instances of a new class of models, namely the Outcome-Driven Process Reward Model. This model class bridges the gap between outcome-based evaluation and process-based supervision by dynamically connecting intermediate reasoning steps to terminal outcomes, enabling more granular feedback during training without the need.
Additionally, the blueprint’s extensive set of operators can inspire the development of innovative reasoning strategies, such as advanced tree-based searches, multi-step refinement processes, or hybrid search algorithms that adapt dynamically to the task’s complexity. These strategies can be tailored using the token probability distribution analysis tools provided, leading to more effective generation strategies that optimize reasoning steps through probabilistic insights. The blueprint also provides a foundation for developing nested architectures where reasoning structures such as trees and graphs are embedded hierarchically. These designs can address multi-layered reasoning tasks, expanding the scope of RLM applications to domains requiring deep, structured reasoning processes.
Scalability remains a key focus of this work. The blueprint’s modular design supports future scalable cloud deployments that enable efficient distribution of compute-intensive tasks across cloud infrastructures. These deployments will not only enhance scalability but also optimize cost and resource utilization, making RLMs more accessible for real-world applications.
By exploring and integrating these ideas, this work aims to empower the next generation of reasoning language models, democratize access to advanced reasoning capabilities, and foster innovation across research and industry. The blueprint’s versatility, combined with the x1 platform, will make it one of the factors in the progress in RLM research and applications.
@@ Acknowledgements
We thank Nicolas Dickenmann for writing the initial MCTS codebase. We thank Hussein Harake, Colin McMurtrie, Mark Klein, Angelo Mangili, and the whole CSCS team granting access to the Ault, Piz Daint and Alps machines, and for their excellent technical support. We thank Timo Schneider for help with infrastructure at SPCL. This project received funding from the European Research Council (Project PSAP, No. 101002047), and the European High-Performance Computing Joint Undertaking (JU) under grant agreement No. 955513 (MAELSTROM). This project received funding from the European Union’s HE research and innovation programme under the grant agreement No. 101070141 (Project GLACIATION). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017103.
@@ Appendix A Mathematical Foundation of Markov Decision Processes for Reasoning Tasks
y so far to put this on the first page of the appendix
TABLE II: Overview of mathematical notation used in the paper
@@
$r(s)$ The reward received when arriving in state $s$.
@@
In this section, we provide a rigorous mathematical framework for RLMs. We achieve this by integrating the theory of Markov Decision Processes (MDPs) with the Monte Carlo Tree Search (MCTS) algorithm. The MDP serves as a foundational formulation for modeling various types of processes, and it can be applied to model reasoning chains, which constitute the reasoning structure of the RLMs. Simultaneously, MCTS serves as an efficient search algorithm for exploring and navigating the extensive space of possible reasoning chains. The resulting state space is then used as a basis for modeling the RLM. An overview of the notation used in this section is provided in Table II.
@@ A.1 Markov Decision Process
A.1.1 Solving an MDP
Before stating what it means formally to solve an MDP, we first need several definitions.
A trajectory $\tau_{\pi}=(s_{0},a_{0},\ldots,s_{T},a_{T},s_{T+1})$ is a sequence of interleaved states and actions, selected according to the policy $\pi$ (see below for the policy definition). Each trajectory starts at an initial state $s_{0}\in\mathcal{S}$ and ends with $s_{T+1}\in\mathcal{S}$ which represents the terminal state where no further actions can be taken.
A policy $\pi(s)$ is a function assigning a probability distribution over the action space to a given state $s$; $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ where $\Delta(\mathcal{A})$ is a set of probability distributions over action space $\mathcal{A}$. The expression $\pi(a\mid s)$ denotes the probability of selecting the action $a$ in the state $s$ according to the policy $\pi$.
State value function $V_{\pi}(s_{t})$ represents the expected cumulative future reward for a given state $s_{t}$ under policy $\pi$:
@@ A.1.1 Solving an MDP
where $T$ is a predefined time-horizon. Note that, in order to obtain the state $s_{k+1}$, an action $a_{k}$ is first derived by sampling from a distribution $\pi(s_{k})$. Once the action $a_{k}$ is chosen, the environment dynamics $p(s_{k+1}\mid s_{k},a_{k})$ determine the probability distribution of the next state $s_{k+1}$.
@@
The State-Action value function $Q(s_{t},a_{t})$ Oftentimes, it is useful to use the state-action value function $Q(s_{t},a_{t})$ instead of the state value function. Specifically, the state-action value function $Q(s_{t},a_{t})$ extends the state value function so that the function value is defined on a state and a specific action $a_{t}$:
@@
$\stackrel{{\scriptstyle}}{{=}}r(s_{t},a_{t})+\gamma\mathbb{E}_{s_{t+1}}\left[V_{\pi}(s_{t+1})\mid s_{t},a_{t}\right],$
where Bellman’s equation is used in the second equality.
A.1.2 MDPs in the RLM Setting
In the context of RLMs, a state $s\in\mathcal{S}$ is typically defined as a sequence of reasoning steps $s=(z_{0}\dots z_{n})$, where each reasoning step $z_{i}$ is a sequence of $M_{i}$ tokens $z_{i}=(t_{i}^{0},\dots,t_{i}^{M_{i}})$. Each $t_{i}^{j}$ is a token from the RLM’s vocabulary, and the total number of tokens per reasoning step $M_{i}$ can vary. One can use a special token $t_{M_{i}}=t_{end}$ to indicate the end of the reasoning step. Typically, the initial query $q$ is used as the first reasoning step $z_{0}=q$. In the study of RLMs, an action $a\in\mathcal{A}_{s}$ usually represents appending a new reasoning step $z^{(a)}$ to the current state $s=(z_{0},...,z_{n})$ resulting in a new state $s^{\prime}=\left(z_{0},...,z_{n},z^{(a)}\right)$. Since every action $a$ is uniquely associated with exactly one reasoning step $z^{(a)}$ for every $s=(z_{0},...,z_{n})$ and $s^{\prime}=(z_{0},...,z_{n},z_{n+1})$, we have
$p(s,a,s^{\prime})=\begin{cases}1&\text{ if }z_{n+1}=z^{(a)}\\ 0&\text{ if }z_{n+1}\neq z^{(a)}\\ \end{cases}$
The definition of the reward function depends on the specific task. A reward commonly seen in reasoning tasks assigns non-zero reward only in the terminal states and hence only at the final reasoning step. This approach reflects the fact that for most tasks, the only final answer can be evaluated against the ground-truth solution to the original query. We call such reward functions sparse to clearly distinguish it from other setting in which intermediate rewards can be observed by the algorithm in the non-terminal states. The discount factor $\gamma$ determines how future rewards influence the current decision-making process. A higher discount factor ($\gamma\rightarrow 1$) places greater emphasis on long-term reasoning success, allowing the model to generate long reasoning sequences, while a lower discount factor prioritizes immediate rewards, incentivizing faster progress and shorter reasoning sequences.